how much more would you need to deposit at end of every month to accumulate $287500 over 18 years if the interest rate earned is 4.5% compounded annually instead of 4.5% compounded monthly?
Answer #1
FVOrdinary Annuity = C*(((1 + i/(f*100))^(n*f) -1)/(i/(f*100)))
C = Cash flow per period
i = interest rate
n = number of payments I f = frequency of payment
287500= Cash Flow*(((1+ 4.5/1200)^(18*12)-1)/(4.5/1200))
Cash Flow = 866.31
Using Calculator: press buttons "2ND"+"FV" then assign
FV =287500
I/Y =4.5/12
N =18*12
PV = 0
CPT PMT
Using Excel
=PMT(rate,nper,pv,fv,type)
=PMT(4.5/(12*100),12*18,,287500,)
EAR = [(1 +stated rate/no. of compounding periods) ^no. of compounding periods - 1]* 100
? = ((1+4.5/(12*100))^12-1)*100
Effective Annual Rate% = 4.594
FVOrdinary Annuity = C*(((1 + i/(f*100))^(n*f) -1)/(i/(f*100)))
C = Cash flow per period
i = interest rate
n = number of payments I f = frequency of payment
287500= Cash Flow*(((1+ 4.594/1200)^(18*12)-1)/(4.594/1200))
Cash Flow = 858.1
Using Calculator: press buttons "2ND"+"FV" then assign
FV =287500
I/Y =4.594/12
N =18*12
PV = 0
CPT PMT
Using Excel
=PMT(rate,nper,pv,fv,type)
=PMT(4.594/(12*100),12*18,,287500,)

Difference = 866.31-858.1

= 8.21